OpenStudy (anonymous):

Simplify (6)/(k)/(y)+2

OpenStudy (anonymous):

\[\large \frac{\frac{6}{k}}{y+2}\]??

OpenStudy (anonymous):

its 6 over k over y+2

OpenStudy (anonymous):

\[\large \frac{6}{\frac{k}{y+2}}\]

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

to divide by a fraction, multiply by the reciprocal, just like with numbers \[6\times \frac{y+2}{k}=\frac{6(y+2)}{k}\]

OpenStudy (anonymous):

thanks